PRELIMINARY OBSERVATIONS OF THE A-X SYSTEM OF THE CALCIUM MONOMETHOXIDE RAD-ICAL : THE ORIGIN BAND

P. CROZET, F. MARTIN, \underline{A. J. ROSS}, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment A. Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France; A. G. ADAM and M. DICK, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2.

Laser excitation spectra of the A ^{2}E ← X ^{2}A_{1} system of calcium methoxide were recorded at the University of New Brunswick. CaOCH_{3} radicals were produced in a laser ablation source, using a 1.2 % mixture of methanol in He as a precursor. High resolution spectra were recorded in the range 15870-15985 cm^{−1} (F.W.H.M. 250 MHz) using a tunable, single-mode cw dye laser (CR 699 + autoscan), operating with DCM dye. The spectra exhibit bands assigned to the two spin-orbit components of the origin band A ^{2}E_{3/2} ← X ^{2}A_{1} and A ^{2}E_{1/2} ← X ^{2}A_{1} . Resolved J and K structures are observed, and tentatively assigned by comparison with synthetic spectra generated from a Hamiltonian given in the literature^{a} and spectroscopic data^{b}^{c} available for the electronic ground state.

aX. Liu, C.P. Damo, T.-Y. D. Liu, S. C. Foster, P. Misra, L. Yu and T. A. Miller J. Phys. Chem \underline{93} 2266-2275 (1989)

bC. J. Whitham, S. A. Beaton, Y.Ito and J. M. Brown J. Mol. Spectrosc. \underline{191} 286-294 (1998)

cK. C. Namiki, J. S. Robinson and T. C. Steimle J. Chem. Phys. \underline{13} 5283-5289 (1998)